OpenStudy (anonymous):

Solve the absolute value equation. Check your solutions. | 5x + 13| = –7

OpenStudy (anonymous):

hint.. from: \[\LARGE |5x+13|=-7\] we have two equation... \[\LARGE 5x+13=-7\] and \[\LARGE 5x+13=7\] solve them for X you'll have your answer :)

OpenStudy (anonymous):

remember the two cases ? same principle applies here.